OpenStudy (anonymous):

Will give medal! please help! what is the inverse of the given relation y=7x^2-3

OpenStudy (jdoe0001):

\(\bf \color{red}{y}=7x^2-3\qquad inverse\implies x=7\color{red}{y}^2-3\) notice, all we did was swap the variables about then solve for "y"

OpenStudy (anonymous):

ok yes i understand so far

OpenStudy (jdoe0001):

so if you solve for "y", you'd get the inverse

OpenStudy (anonymous):

would it be x=sqrt( y+3/7 ?

OpenStudy (jdoe0001):

well.. you're meant to solve for "y"... the above shows a solution for "x"

OpenStudy (anonymous):

oh sorry, like y^2= x+3/7?

OpenStudy (jdoe0001):

.... wait.. a second

OpenStudy (jdoe0001):

well... if we.... ahemm... follow that EXPRESSION, then we'd end up with \(\bf x=7y^2-3\implies \cfrac{x+3}{7}=y^2\implies \sqrt{\cfrac{x+3}{7}}=y\) however, notice the graph of your original function --> http://fooplot.com/#W3sidHlwZSI6MCwiZXEiOiI3eF4yLTMiLCJjb2xvciI6IiNGMjE4MTgifSx7InR5cGUiOjEwMDB9XQ-- notice anything?

OpenStudy (anonymous):

no im confused

OpenStudy (jdoe0001):

well... is a parabola.. so.... the definition of an "inverse function" is obtained by the procedure we just did, yes, however, ONLY ONE-TO-ONE functions have an inverse function, and a parabola isn't a ONE-TO-ONE function, so by definition the original function, which is just a parabola, wouldn't have an inverse function

OpenStudy (jdoe0001):

a one-to-one function has to pass the "vertical line test" as well as the "horizontal line test", a parabola passes the "vertical" one, not the "horizontal" one though, thus

OpenStudy (anonymous):

umm , so your saying their is no inverse?

OpenStudy (jdoe0001):

yes, because the original function is not a one-to-one function, it's a function nonetheless, just not a one-to-one, and by definition that's required for an inverse function to be

OpenStudy (anonymous):

well their are answer options so their has to be one

OpenStudy (jdoe0001):

well, the expression we'd get would be \(\bf x=7y^2-3\implies \cfrac{x+3}{7}=y^2\implies \sqrt{\cfrac{x+3}{7}}=y\)

OpenStudy (anonymous):

here:

OpenStudy (jdoe0001):

ohhh smokes... .I see the semantics.... ok... it says RELATION, no function.... ok then my bad

OpenStudy (anonymous):

ohh ok

OpenStudy (jdoe0001):

well, a root has a \(\bf \pm \) results, so you can always write that as \(\bf x=7y^2-3\implies \cfrac{x+3}{7}=y^2\implies \pm\sqrt{\cfrac{x+3}{7}}=y\)

OpenStudy (anonymous):

ok yes, thank u!

OpenStudy (jdoe0001):

yw